Defense spending ~ The US federal budget for 2020 allocated $705 billion for defense and military costs. A 2020 Gallup poll concluded that the proportion of Americans, who believe that the defense budget allocation is appropriate, is 0.5047. A sociologist at MSU wants to determine if the actual proportion of Americans who believe that the defense budget is appropriate is different from the value reported by the Gallup poll.

A normal distribution is used to carry out the following hypothesis test:

HO : P=.5047 vs HA p =/ .5047

The test statistic for this hypothesis test is 1.2775. Find the p-value. Give your answer to 4 decimal places.
Answer #1

Solution :

This is the two tailed test .

The null and alternative hypothesis is

H0 : p =0.5047

Ha : p 0.5047

P(z >1.2775 ) = 1 - P(z <1.2775 ) = 1-0.8993=0.1007*2=0.2014

P-value = 0.2014 (using z table)